OpenStudy (anonymous):

which answer describes the solution to each equation? column a 1. x-6/2 2. 2+6x-7=8+2x-10+4x 3. 3x+5=1/5(15 x+10) 4. 1/4(16+8x)=3x-1-x+5 column b a. one solution b. no solution c. an infinite number of solutions

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

1. b 2. c 3 c and 4. a?

OpenStudy (anonymous):

@iGreen. @AnswerMyQuestions

OpenStudy (anonymous):

Um, the columns are messed up, can you fix them?

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

sorry it took forever to type then os messed up so i screen shoted it.

OpenStudy (anonymous):

oops i forgot part of 1.

OpenStudy (anonymous):

it is x-6/2 = x+13

OpenStudy (anonymous):

Oh, okay, first we add 6/2 to both sides. What's 13 + 6/2?

OpenStudy (anonymous):

6/2 is 3.. So 13 + 3 = ?

OpenStudy (anonymous):

@Demonx341

OpenStudy (anonymous):

16

OpenStudy (anonymous):

Yes, now we have: x = x + 16 Now subtract 'x' to both sides, what's x - x?

OpenStudy (anonymous):

umm x?

OpenStudy (anonymous):

or 0?>

OpenStudy (anonymous):

Yes, 0..so this one has 0 solutions.

OpenStudy (anonymous):

2. a?

OpenStudy (anonymous):

No, it's actually the same as the other one..it has 0 solutions.

OpenStudy (anonymous):

\[\color{red}{Is}\color{green}{\\ A}\color{pink}{\ \\ B?}\]

OpenStudy (anonymous):

xD not a 2.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Is 2. B?

OpenStudy (anonymous):

Yes, that's correct.

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

3. Hmmm

OpenStudy (anonymous):

3. A 4. C?